OpenStudy (anonymous):

How much would $200 invested at 4% interest compounded monthly be worth after 8 years? Round your answer to the nearest cent. a(t)=p(1+r/2)^nt A. $205.40 B. $273.71 C. $322.83 D. $275.28

OpenStudy (danjs):

r = 4% = 0.04 compounded monthly = 12 times per year, for t = 8 years

OpenStudy (danjs):

principle P = 200

OpenStudy (danjs):

If you put those values in you get \[a(8) = 200*(1+\frac{ 0.04 }{ 12 })^{12*8}\]

OpenStudy (anonymous):

Would it be A?

OpenStudy (danjs):

hmm

OpenStudy (danjs):

1+4/1200 = 301/300 \[a(8) = 200*(\frac{ 301 }{ 300 })^{96}\] with a calculator i got a(8) is about 275.279